# EDITOR: Effective and Interpretable Prompt Inversion for Text-to-Image Diffusion Models

### v1

@@ Abstract
Text-to-image generation models (e.g., Stable Diffusion) have achieved significant advancements, enabling the creation of high-quality and realistic images based on textual descriptions. Prompt inversion, the task of identifying the textual prompt used to generate a specific artifact, holds significant potential for applications including data attribution, model provenance, and watermarking validation. Recent studies introduced a delayed projection scheme to optimize for prompts representative of the vocabulary space, though challenges in semantic fluency and efficiency remain. Advanced image captioning models or visual large language models can generate highly interpretable prompts, but they often lack in image similarity. In this paper, we propose a prompt inversion technique called EDITOR for text-to-image diffusion models, which includes initializing embeddings using a pre-trained image captioning model, refining them through reverse-engineering in the latent space, and converting them to texts using an embedding-to-text model. Our experiments on the widely-used datasets, such as MS COCO, LAION, and Flickr, show that our method outperforms existing methods in terms of image similarity, textual alignment, prompt interpretability and generalizability. We further illustrate the application of our generated prompts in tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
@@ 1 Introduction
Text-to-image generation models have made remarkable progress, greatly enhancing the creation of visual content. Advanced deep learning techniques enable models like Parti [1], DALL-E [2], and Stable Diffusion [3] to produce high-quality, realistic, and creative images based on textual descriptions. A key factor in successfully generating high-quality images is the design of prompts. Well-crafted prompts guide these models to generate images that match the desired details and styles. As such, prompts used to generate high-quality images have become a critical component of the ecosystem and valuable IP assets [4, 5, 6].
Prompt inversion is the task of reconstructing the textual prompt used to generate a specific image. It has various applications in trustworthy artificial intelligence (AI), such as data attribution, model provenance, and watermarking validation [7, 8, 9, 10, 6]. However, as summarized in Table 1, current methods face two main challenges: they either lack prompt interpretability or fall short in image similarity.
@@
C1. Prompt Interpretability? Existing prompt inversion methods [11, 12] are localized in the latent space and optimize for prompts that are representative of the vocabulary space. Specifically, As shown in 2(a), they optimize the latent embeddings without processing them through transformer layers and project them back to the nearest neighbors in the vocabulary space. As a result, the inverted prompts often lack semantic fluency and these methods are computationally inefficient, searching within discrete spaces. Prompts that are difficult to interpret also severely hinder their generalizability, greatly reducing their value.
C2. Image Similarity? A possible solution to get interpretable prompts involves directly employing advanced image captioning models or visual large language models to obtain accurate and fluent prompts. For example, BLIP-2 [13], LLaVA-1.5 [14] and Janus-Pro [15]. However, this direct generation often lacks similarity in image generation, especially when the goal is to produce a description that is maximally aligned with a specific image using a particular difuusion model.
@@
Figure 2: Existing works optimize embedding before transformer layer and project them to nearest token embeddings, often producing incoherent results; EDITOR optimizes embedding after transformer layer and converts the final optimized embeddings into prompts. The red arrow indicates the backward pass in gradient-based optimization.
In contrast, we propose a prompt inversion method, called EDITOR, for text-to-image diffusion models to effectively generate prompts that are both accurate and elegant. The overall process starts by initializing the latent embedding using a pre-trained image captioning model. Then, we optimize the prompts to recover a more accurate and contextually appropriate prompt. In particular, as shown in 2(b), we first initialize the latent text embedding, process it through transformer layers to obtain the contextualized embedding, then apply gradient-based optimization to obtain the reconstructed embedding, and finally invert it back to the text space using a trained embedding-to-text model [16]. These three steps work together to ensure that the inverted prompt maintains coherence and interpretability. Figure 1 illustrates examples of our prompt inversion result and comparison with existing work [12, 11]. Compared to baselines, our method produces much more coherent and interpretable prompts. Furthermore, the reverse-engineering step ensures that the generated prompt remains semantically aligned with the target image, offering improved image quality compared to directly using an image captioning model.
@@
Our results on the widely-used datasets (e.g., MS COCO [17], LAION [18], and Flickr [19]) under state-of-the-art text-to-image generation model (e.g., Stable Diffusion [3]) show that EDITOR is more effective in both image similarity and prompt quality than existing methods. Our contributions are summarized as follows:
@@
  • We introduce a novel technique called EDITOR for inverting prompts in text-to-image diffusion models. Unlike existing methods that use discrete optimization with projection on vocabulary, EDITOR improves optimization effectiveness and efficiency through proper initialization and reverse-engineering in a continuous latent space.
@@
  • The EDITOR employs a novel three-step pipeline including initialization, revere-engineering and embedding inversion. This framwork surpasses existing prompt inversion methods like PEZ [12] and PH2P [11] in prompt interpretability and outperforms image captioning models in image similarity.
@@
  • We demonstrate that EDITOR produces semantically meaningful prompts excelling in image similarity, text alignment, and interpretability across three standard datasets. Moreover, they enable flexible downstream tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
@@ 2 Related Work
Text-to-Image Generation. Early approaches to text-to-image generation were dominated by generative adversarial networks (GANs) [20, 21] to generate style-controllable images. Inspired by the success of autoregressive transformer [22], recent methods [23, 24] generate images sequentially by modeling the conditional probability of pixels or patches.
Significant progress in high-resolution image synthesis has been made with diffusion models [25] that gradually refines an initial noisy input into a coherent image. Conditional diffusion models [2, 26] leverage text encoders [27, 28] to map texts into latent embeddings for controllable generation. To improve efficiency, latent diffusion models (LDMs) [3, 29, 30], perform the diffusion process in a compressed latent space instead of the pixel space. Given an image $x$, an encoder $\mathcal{E}(\cdot)$ compresses the image into a latent representation $z$, and a decoder $\mathcal{D}(\cdot)$ reconstructs the image back into pixel space, such that $x\approx\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$. Here, $z$ represents the latent representations of the image. During the diffusion process, noise is incrementally added to $z$, resulting in a sequence of latent representations $z_{t}$ for each timestep $t\in\{0,1,\dots,T\}$, where $T$ is the total number of timesteps. A U-Net [31] $\epsilon_{\theta}(z_{t},t)$ is trained as a denoising module to predict the noise component within $z_{t}$. In conditional text-to-image diffusion models, the denoising process is guided by conditional text inputs. A transformer-based text encoder $\mathcal{T}(\cdot)$ is employed to convert a given prompt $p$ into a corresponding latent text embedding $c:=\mathcal{T}(p)$. The conditional U-Net $\epsilon_{\theta}$ takes $(z_{t},t,c)$ as input, and the training objective of the conditional latent diffusion model is formalized as
@@
Inversion for Diffusion Models. Previous inversion techniques in diffusion models aim to enhance controllability and personalization in text-to-image generation by embedding visual concepts into the model’s textual or latent space [32, 33, 34, 35, 36]. DreamBooth [34] fine-tunes diffusion models using specific reference images to learn new visual features for personalized image generation. Textual inversion [32] encodes visual concepts as unique tokens in the model’s vocabulary, enabling customized image generation through textual prompts. [36] presents an encoder-based approach for fast personalization of text-to-image models. Similarly, ELITE [35] extends textual inversion by encoding multiple visual concepts into the textual embedding space, allowing the combination of various concepts within a single prompt. Moreover, [33] extends prompt conditioning, where different embeddings are injected into different layers of the U-net. Although these methods enhance personalization and controllability in text-to-image generation, they often lack interpretability and generalization, limiting their applicability to diverse contexts. One solution is to invert image features into natural language representations. Inspired by discrete optimization, recent works focus on optimizing textual prompts to better align with image features [12, 11][12] introduces a gradient-based discrete optimization method for hard prompt discovery with CLIP [37] and [11] proposes hard prompt optimization that aligns prompts with image features directly in diffusion models. These approaches bridge the gap between visual and textual representations, allowing for more intuitive manipulation of generated images through natural language. However, these inverted prompts often lack semantic fluency and naturalness, resulting in outputs unintelligible to humans. In contrast, our method ensures image similarity while making prompts more human-understandable, thereby enhancing their interpretability and generalizability.
@@ 3 Methodology
3.1 Problem Formulation
We aim to invert prompts from text-to-image generation models. Following existing works [12, 11], we formulate the problem as follows:
@@ 3.1 Problem Formulation
Engineer’s Capability. Same as [11], we assume that we have white-box access to a latent diffusion model $\mathcal{M}$, allowing them to access intermediate outputs and compute the model’s gradients. In addition, the attacker has access to other public models and tools, which aids in more efficient prompt optimization.
3.2 Overview
@@
Input:   Text-to-Image Diffusion Model: $\mathcal{M}$, Image Captioning Model: $\mathcal{M}_{c}$, Embedding-to-Text Model: $\mathcal{M}_{t}$, Target Image: $\bm{x}$.
@@
7:     for $e\leq\rm{max\_epoch}$ do
8:         $loss=\mathcal{L}\left(\mathcal{M}(\bm{c}),\bm{x}\right)$
9:         $\Delta_{\bm{c}}=\frac{\partial loss}{\partial\bm{c}}$
10:         $\bm{c}=\bm{c}-lr\cdot\Delta_{\bm{c}}$
@@
The overview of EDITOR is shown in Figure 3, consisting of three main steps: initialization, reverse-engineering, and embedding inversion. Firstly, we initialize the latent text embedding using an image captioning model, which provides a starting point with a semantically meaningful prompt, addressing the issue of isolated and incoherent tokens. Secondly, we apply reverse-engineering to this initial embedding to obtain a reconstructed embedding. Finally, we convert the reconstructed embedding back to text using an embedding-to-text model.
Algorithm. The overall gradient-based optimization algorithm is shown in Algorithm 1. We begin by using an image captioning model to transform the target image into a textual description, which is then embedded by the text encoder of the diffusion model. Next, we iteratively refine this embedding by calculating the gradient of a reconstruction loss that compares the diffusion model’s generated output against the original image, and we update the embedding accordingly over multiple epochs. Finally, we feed the refined embedding to a text-to-image model, thereby recovering an "inverted prompt" that closely aligns with the visual content of the target image.
3.3 Initialization for Latent Text Embedding
We introduce a novel initialization strategy that leverages pre-trained image captioning models to generate an initial prompt for the target image. This approach ensures that the inverted prompt begins with a semantically meaningful and contextually relevant prompt, thereby reducing the search space and aligning the embedding more closely with the target distribution. Specifically, for a given image $x$, we utilize an image captioning model to generate an initial prompt $p_{0}$. This initial prompt is then encoded into a starting point fot optimization using the text encoder $\mathcal{T}$, which maps the text into the latent space of the diffusion model.
It is worth noting that the initialization strategy is unique to our method, benefiting from the optimization in a continuous latent space. Unlike hard prompt optimization methods such as PEZ [12] and PH2P [11], our approach does not significantly disrupt the syntactic structure of the initial prompt. In contrast, as shown in section 6, even with initialization, PEZ and PH2P produce token combinations with high perplexity.
3.4 Reverse-engineering for Latent Text Embedding
With the initial prompt, we reconstruct latent text embeddings using input reverse-engineering. Given a target image $x$ and a latent diffusion model, the objective is to find an embedding $c^{*}$ such that the generated image $\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(c^{*},n))$ minimizes the distance to $x$. This input reverse-engineering task is formulated as:
@@ 3.4 Reverse-engineering for Latent Text Embedding
where $\mathcal{L}$ is a distance metric that measures the similarity between two images, referred to as the reconstruction loss. The latent random noise $n$ is fixed, and the parameters of the latent diffusion model are held constant.
@@
In contrast to existing methods [12, 11] that optimize the embeddings before passing through the transformer layers and then project them onto the corresponding vocabulary, we directly optimize the output of the text encoder in the latent space. This approach offers two key advantages: (1) it improves optimization efficiency by bypassing the projection step, which can be computationally expensive and slow, and it ensures better semantic consistency and coherence. As shown in  Table 14, EDITOR improves optimization efficiency by more than double compared to PH2P; (2) by optimizing the text encoder’s output, which has already undergone deeper semantic processing, the generated text remains more natural, contextually relevant, and grammatically correct, while also avoiding the disjointed or redundant word choices that can arise in discrete optimization.
3.5 Embedding Inversion
It is also challenging to get text prompts from the reconstructed embedding, as the embedding is continuous and highly contextualized, making it impossible to be projected to the nearest token in the vocabulary, as is done in PEZ [12] and PH2P [11]. To solve this problem, we leverage an embedding-to-text model [16] to convert the reconstructed embedding back into text prompts. Let $c\in\mathbb{R}^{d}$ be a target (latent) embedding. We first train a zero-step model $M_{\mathrm{zero}}$ that directly maps $c$ to text $x$, estimating the conditional probability $p_{\mathrm{zero}}(x\mid c)$. This training uses a maximum likelihood objective:
$\mathcal{L}_{\text{ML}}=-\sum{(x,c)}\log p_{\mathrm{zero}}\bigl{(}x\mid c\bigr{)}.$ (4)
Once trained, $M_{\mathrm{zero}}$ generates an initial hypothesis $\hat{x}$ by sampling from $p_{\mathrm{zero}}(x\mid c)$. Next, we introduce a correction model $M_{\mathrm{corr}}$ to iteratively refine $\hat{x}$. In each refinement step $k$, the model produces a revised text $x^{(k)}$ conditioned on 1) the target embedding $c$, 2) the current hypothesis $x^{(k-1)}$, and 3) the embedding $\mathcal{T}\bigl{(}x^{(k-1)}\bigr{)}$. Formally:
$p_{\mathrm{corr}}\bigl{(}x^{(k)}\mid c;x^{(k-1)};\mathcal{T}(x^{(k-1)})\bigr{)}.$ (5)
@@ 3.5 Embedding Inversion
$\mathcal{L}_{\text{CE}}=-\sum{(x^{(k)},c)}\log p_{\mathrm{corr}}(x^{(k)}|c).$ (6)
During inference, beam search is applied to generate multiple refined hypotheses, and the final output is selected to minimize the distance $\bigl{\|}\,\mathcal{T}(x)\,-\,c\,\bigr{\|}$ with respect to the ground-truth embedding $c$.
Compared to PEZ [12] and PH2P [11], our embedding-to-text process significantly improves the interpretability of the inverted prompt. Due to the attention mechanism, the embedding at each position can incorporate information from previous tokens, making it more likely that phrases aligned with the image’s semantics are preserved in the generated text, rather than just individual words. Additionally, when the optimized embedding introduces new or uninterpretable features, the highly interpretable training data enables the model to convert these features into more understandable and coherent text.
Table 2: Image similarity comparison with baselines.
Dataset Metric PEZ PH2P EDITOR
MS COCO CLIP Similarity $\uparrow$ 0.745 0.789 0.796
LPIPS Similarity $\downarrow$ 0.477 0.467 0.414
LAION CLIP Similarity $\uparrow$ 0.756 0.826 0.826
LPIPS Similarity $\downarrow$ 0.455 0.427 0.401
Flickr CLIP Similarity $\uparrow$ 0.722 0.755 0.776
LPIPS Similarity $\downarrow$ 0.442 0.439 0.424
Table 3: Prompt quality comparison with baselines.
Dataset Metric PEZ PH2P EDITOR
MS COCO Precision $\uparrow$ 0.791 0.801 0.900
@@
[…]
PPL $\downarrow$ 8,837.797 11,078.994 80.659
LAION Precision $\uparrow$ 0.783 0.794 0.841
[…]
PPL $\downarrow$ 8040.401 9621.671 187.299
Flickr Precision $\uparrow$ 0.774 0.798 0.892
[…]
PPL $\downarrow$ 9,425.353 11,804.100 90.320

### v2

@@ Abstract
Text-to-image generation models (e.g., Stable Diffusion) have achieved significant advancements, enabling the creation of high-quality and realistic images based on textual descriptions. Prompt inversion, the task of identifying the textual prompt used to generate a specific artifact, holds significant potential for applications including data attribution, model provenance, and watermarking validation. Recent studies introduced a delayed projection scheme to optimize for prompts representative of the vocabulary space, though challenges in semantic fluency and efficiency remain. Advanced image captioning models or visual language models can generate highly interpretable prompts, but they often lack in image similarity. In this paper, we propose a prompt inversion technique called EDITOR for text-to-image diffusion models, which includes initializing embeddings using a pre-trained image captioning model, refining them through reverse-engineering in the latent space, and converting them to texts using an embedding-to-text model. Our experiments on the widely-used datasets, such as MS COCO, LAION, and Flickr, show that our method outperforms existing methods in terms of image similarity, textual alignment, prompt interpretability and generalizability. We further illustrate the application of our generated prompts in tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
@@ 1 Introduction
Text-to-image generation models have made remarkable progress, greatly enhancing the creation of visual content. Advanced deep learning techniques enable models like Parti (Yu et al., 2022), DALL-E (Ramesh et al., 2022), and Stable Diffusion (Rombach et al., 2022) to produce high-quality, realistic, and creative images based on textual descriptions. A key factor in successfully generating high-quality images is the design of prompts. Well-crafted prompts guide these models to generate images that match the desired details and styles. As such, prompts used to generate high-quality images have become a critical component of the ecosystem and valuable IP assets (Wang et al., 2024a; Zhang et al., 2023; Shen et al., 2024).
Prompt inversion is the task of reconstructing the textual prompt used to generate a specific image. It has various applications in trustworthy artificial intelligence (AI), such as data attribution, model provenance, and watermarking validation (Wang et al., 2024b; c; 2023; Naseh et al., 2024; Shen et al., 2024). However, as summarized in Table 1, current methods face two main challenges: they either lack image similarity or fall short in prompt interpretability.
@@
A straightforward way to obtain interpretable prompts is to apply captioning models or advanced vision language models, such as BLIP-2 (Li et al., 2023), LLaVA-1.5 (Liu et al., 2024), or Janus-Pro (Chen et al., 2025). While these models generate fluent descriptions, they fail to ensure high similarity when re-generating the image with a diffusion model. On the other hand, optimization-based approaches such as PEZ (Wen et al., 2024) and PH2P (Mahajan et al., 2024) update latent embeddings and repeatedly project them onto the vocabulary. However, these discrete projections disrupt semantic continuity, causing unreadable prompts, computational inefficiency. More importantly, the projection step incurs severe embedding discrepancy, with empirical results indicating that the cosine similarity between the optimized embedding and its projected counterpart often falls to only 0.167. Such a drastic degradation implies that even if optimization successfully approaches a near-optimal embedding, the subsequent projection may push the representation far away from this solution, deviating negatively from the optimum. Consequently, the optimization process becomes unstable and less efficient, limiting the effectiveness of these methods in recovering high-quality prompts.
@@
Figure 2: Existing works optimize the token embedding before transformer layer and project them to nearest token embeddings, often producing incoherent results; EDITOR optimizes the contextual embedding after transformer layer and converts the final optimized embeddings into prompts. The red arrow indicates the backward pass in gradient-based optimization.
In this paper, we propose a new prompt inversion method, called EDITOR, for text-to-image diffusion models to effectively generate prompts that are both accurate and elegant. Instead of projecting the embeddings back to the prompts at every iteration as done by existing approaches (Wen et al., 2024; Mahajan et al., 2024), EDITOR directly optimizes the representation in the continuous space (e.g., contextual embeddings). This ensures that the optimized embeddings are optimal and can induce the intended image. To obtain discrete prompts, EDITOR leverages an embedding-to-text model (Morris et al., 2023) for projection. Unlike prior work, which simply projects embeddings to their nearest prompts—causing substantial discrepancies—we utilize text-representation pairs collected from the subject diffusion model to train our embedding-to-text model. Additionally, EDITOR is equipped with a correction model that iteratively refines prompts, pulling their embeddings closer to their originally optimized counterparts. This design ensures that the optimized representation can be projected into an in-distribution space and is therefore more semantically aligned with the original prompt. Figure 2 illustrates the comparison between existing approaches and our method EDITOR in optimizing embeddings and obtaining prompts. Figure 1 presents examples of our prompt inversion result and comparison with PH2P (Mahajan et al., 2024). Compared to the baselines, our method produces much more coherent and interpretable prompts. Furthermore, the reverse-engineering step ensures that the generated prompt remains semantically aligned with the target image, offering improved image quality compared to directly using an image captioning model.
@@
Our results on the widely-used datasets (e.g., MS COCO (Lin et al., 2014), LAION (Schuhmann et al., 2022), and Flickr (Young et al., 2014)) under state-of-the-art text-to-image generation model (e.g., Stable Diffusion (Rombach et al., 2022)) show that EDITOR is more effective in both image similarity and prompt quality than existing methods. Our contributions are summarized as follows:
@@
  • We introduce a novel technique called EDITOR for inverting prompts in text-to-image diffusion models. Unlike existing methods that rely on discrete optimization with token embedding projection onto the vocabulary—a process that disrupts semantic continuity and causes severe embedding discrepancy—EDITOR improves optimization effectiveness and efficiency through contextual embedding optimization in a continuous latent space.
@@
  • EDITOR employs a novel three-step prompt inversion pipeline including initialization, revere-engineering and embedding inversion. This framwork surpasses existing prompt inversion methods like PEZ (Wen et al., 2024) , PH2P (Mahajan et al., 2024) and VGD (Kim et al., 2025) in prompt interpretability and outperforms advanced image captioning models and vision language models in image similarity.
@@
  • We demonstrate that EDITOR produces semantically meaningful prompts excelling in image similarity, text alignment, and interpretability across three standard datasets. In addition, EDITOR also achieves strong performance when applied to advanced multi-encoder diffusion models such as SDXL-Turbo and Stable Diffusion 3.5 Medium, showing its robustness across different architectures. Moreover, they enable flexible downstream tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
@@ 2 Related Work
Text-to-Image Generation. Early approaches to text-to-image generation were dominated by generative adversarial networks (GANs) (Zhang et al., 2017; Karras, 2019) to generate style-controllable images. Inspired by the success of autoregressive transformer (Vaswani, 2017), recent methods (Ding et al., 2021; Esser et al., 2021) generate images sequentially by modeling the conditional probability of pixels or patches. Significant progress in high-resolution image synthesis has been made with diffusion models (Ho et al., 2020) that gradually refines an initial noisy input into a coherent image. Conditional diffusion models (Ramesh et al., 2022; Saharia et al., 2022) leverage text encoders (Radford et al., 2021b; Raffel et al., 2020b) to map texts into latent embeddings for controllable generation. To improve efficiency, latent diffusion models (LDMs) (Rombach et al., 2022; Gu et al., 2022; Luo et al., 2023), perform the diffusion process in a compressed latent space instead of the pixel space. Given an image $x$, an encoder $\mathcal{E}(\cdot)$ compresses the image into a latent representation $z$, and a decoder $\mathcal{D}(\cdot)$ reconstructs the image back into pixel space, such that $x\approx\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$. Here, $z$ represents the latent representations of the image. During the diffusion process, noise is incrementally added to $z$, resulting in a sequence of latent representations $z_{t}$ for each timestep $t\in\{0,1,\dots,T\}$, where $T$ is the total number of timesteps. A U-Net (Ronneberger et al., 2015) $\epsilon_{\theta}(z_{t},t)$ is trained as a denoising module to predict the noise component within $z_{t}$. In conditional text-to-image diffusion models, the denoising process is guided by conditional text inputs. A transformer-based text encoder $\mathcal{T}(\cdot)$ is employed to convert a given prompt $p$ into a corresponding latent text embedding $c:=\mathcal{T}(p)$. The conditional U-Net $\epsilon_{\theta}$ takes $(z_{t},t,c)$ as input, and the training objective of the conditional latent diffusion model is formalized as
@@
Soft Prompt Inversion for Diffusion Models. Previous inversion techniques in diffusion models aim to enhance controllability and personalization in text-to-image generation by embedding visual concepts into the model’s textual or latent space (Gal et al., 2022; Voynov et al., 2023; Ruiz et al., 2023; Wei et al., 2023; Gal et al., 2023). DreamBooth (Ruiz et al., 2023) fine-tunes diffusion models using specific reference images to learn new visual features for personalized image generation. Textual inversion (Gal et al., 2022) encodes visual concepts as unique tokens in the model’s vocabulary, enabling customized image generation through textual prompts. Gal et al. (2023) presents an encoder-based approach for fast personalization of text-to-image models. Similarly, ELITE (Wei et al., 2023) extends textual inversion by encoding multiple visual concepts into the textual embedding space, allowing the combination of various concepts within a single prompt. Moreover, Voynov et al. (2023) extends prompt conditioning, where different embeddings are injected into different layers of the U-net. Although these methods enhance personalization and controllability in text-to-image generation, they often lack interpretability and generalization, limiting their applicability to diverse contexts.
Hard Prompt Inversion for Diffusion Models. One solution to obtain the hard prompt is to invert image features into natural language representations. Inspired by discrete optimization, recent works (Wen et al., 2024; Mahajan et al., 2024) focus on optimizing textual prompts to better align with image features. Wen et al. (2024) introduces a gradient-based discrete optimization method for hard prompt discovery with CLIP (Mokady et al., 2021) and Mahajan et al. (2024) proposes hard prompt optimization that aligns prompts with image features directly in diffusion models. However, these inverted prompts often lack semantic fluency and naturalness, resulting in outputs unintelligible to humans. Another recent work, Visually Guided Decoding (VGD) (Kim et al., 2025), is gradient-free and uses a large language model plus CLIP-based guidance to generate coherent, human-readable prompts. However, it can generate fluent prompts efficiently but still falls short in terms of achieving high image similarity. In contrast, our method ensures image similarity while making prompts more human-understandable, thereby enhancing their interpretability and generalizability.
@@ 3 Methodology
Problem Formulation. We aim to invert prompts from text-to-image generation models. Following existing works (Wen et al., 2024; Mahajan et al., 2024), we formulate the problem as follows:
@@
  • Engineer’s Capability. Same as Mahajan et al. (2024), we assume that we have white-box access to a latent diffusion model $\mathcal{M}$, allowing them to access intermediate outputs and compute the model’s gradients. In addition, the attacker has access to other public models and tools, which aids in more efficient prompt optimization.
Overview. The overview of EDITOR is shown in Figure 3, consisting of three main steps: initialization, reverse-engineering, and embedding inversion. The overall gradient-based optimization algorithm is shown in Algorithm 1.
@@
Input: Text Decoder $\mathcal{D}$, Generation Process $\mathcal{R}$, U-Net $\epsilon_{\theta}$, Noise $n$, Image Captioning Model $\mathcal{M}_{c}$, E2T Model $\mathcal{M}_{t}$, Target Image $\bm{x}$.
@@
7:  for $e\leq\text{max\_epoch}$ do
8:   $loss=\mathcal{L}\left(\mathcal{D}(\mathcal{R}_{\epsilon_{\theta}}(c,n)),\bm{x}\right)$
9:   $\nabla_{\bm{c}}=\frac{\partial loss}{\partial\bm{c}}$
10:   $\bm{c}=\bm{c}-lr\cdot\nabla_{\bm{c}}$
@@
Figure 4: Comparison of optimization paths. The blue path illustrates continuous optimization. The red path shows optimization with projection in PEZ and PH2P, where each step is displaced due to vocabulary projection.
Initialization for Latent Text Embedding. We leverage pre-trained image captioning models to generate an initial prompt for the target image. This approach ensures that the inverted prompt begins with a semantically meaningful and contextually relevant prompt, thereby reducing the search space and aligning the embedding more closely with the target distribution. Specifically, for a given image $x$, we utilize an image captioning model to generate an initial prompt $p_{0}$. This initial prompt is then encoded into a starting point for optimization using the text encoder $\mathcal{T}$, which maps the text into the latent space of the diffusion model.
It is worth noting that the initialization strategy is unique to our method, benefiting from the optimization in a continuous latent space. Unlike hard prompt optimization methods such as PEZ (Wen et al., 2024) and PH2P (Mahajan et al., 2024), our approach does not significantly disrupt the syntactic structure of the initial prompt. In contrast, as shown in Table 5, even with initialization, PEZ and PH2P produce token combinations with high perplexity.
Reverse-engineering for Latent Text Embedding. With the initial prompt, we reconstruct latent text embeddings using input reverse-engineering. Given a target image $x$ and a latent diffusion model, the objective is to find an embedding $c^{*}$ such that the generated image $\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(c^{*},n))$ minimizes the distance to $x$. This input reverse-engineering task is formulated as:
@@
where $\mathcal{L}$ denotes the mean squared error (MSE) loss, which computes the average squared difference between the generated and target images. The latent random noise $n$ is fixed, and the parameters of the latent diffusion model are held constant.
@@
In contrast to existing methods (Wen et al., 2024; Mahajan et al., 2024) that optimize the token embeddings before passing through the transformer layers and then project them onto the corresponding vocabulary, we directly optimize the output of the text encoder in the latent space. This approach offers two key advantages: (1) it enables more stable convergence by eliminating the repeated projection step. As illustrated in Figure 4, projection often displaces embeddings away from the optimal region, forcing the optimization to restart from a less favorable position. By avoiding this detour, EDITOR maintains the optimization trajectory within the continuous latent space, leading to significantly improved convergence speed and efficiency. (2) by optimizing the text encoder’s output, which has already undergone deeper semantic processing, the generated text remains more natural, contextually relevant, and grammatically correct, while also avoiding the disjointed or redundant word choices that can arise in discrete optimization.
Embedding Inversion. It is also challenging to get text prompts from the reconstructed embedding, as the embedding is continuous and highly contextualized, making it impossible to be projected to the nearest token in the vocabulary, as is done in PEZ (Wen et al., 2024) and PH2P (Mahajan et al., 2024). To solve this problem, we leverage an embedding-to-text (E2T) model (Morris et al., 2023) to convert the reconstructed embedding back into text prompts. Let $c\in\mathbb{R}^{d}$ be a target (latent) embedding. We first train a zero-step model $M_{\mathrm{zero}}$ that directly maps $c$ to text $p$, estimating the conditional probability $P_{\mathrm{zero}}(p\mid c)$. This training uses a maximum likelihood objective:
$\mathcal{L}_{\text{ML}}=-\sum{(p,c)}\log P_{\mathrm{zero}}\bigl(p\mid c\bigr).$ (4)
Once trained, $M_{\mathrm{zero}}$ generates an initial hypothesis $\hat{p}$ by sampling from $P_{\mathrm{zero}}(p\mid c)$. Next, we introduce a correction model $M_{\mathrm{corr}}$ to iteratively refine $\hat{p}$. In each refinement step $k$, the model produces a revised text $p^{(k)}$ conditioned on 1) the target embedding $c$, 2) the current hypothesis $p^{(k-1)}$, and 3) the embedding $\mathcal{T}\bigl(y^{(k-1)}\bigr)$. Formally:
$P_{\mathrm{corr}}\bigl(p^{(k)}\mid c;p^{(k-1)};\mathcal{T}(p^{(k-1)})\bigr).$ (5)
@@
$\mathcal{L}_{\text{CE}}=-\sum{(p^{(k)},c)}\log P_{\mathrm{corr}}(p^{(k)}\mid c).$ (6)
During inference, beam search is applied to generate multiple refined hypotheses, and the final output is selected to minimize the distance $\bigl\|\,\mathcal{T}(p)\,-\,c\,\bigr\|$ with respect to the ground-truth embedding $c$.
We train our embedding-to-text model using text-representation pairs generated by the text encoder of the diffusion model. This design ensures that the optimized representations are mapped back into an in-distribution space, making them more semantically aligned with the original prompts.
Table 2: Vocabulary Projection vs. Embedding Inversion.
Metric Vocabulary Projection Embedding Inversion
Cosine Similarity $\uparrow$ 0.167 0.737
L2 Distance $\downarrow$ 32.934 11.373
Compared to PEZ (Wen et al., 2024) and PH2P (Mahajan et al., 2024), our embedding-to-text process also achieves far lower embedding discrepancy. As shown in Table 2, projection onto the vocabulary space incurs severe embedding discrepancy (cosine similarity of only 0.1671 and L2 distance of 32.93), whereas our embedding inversion achieves markedly better alignment (cosine similarity of 0.7370 and L2 distance of 11.37). This quantitative gap indicates that direct projection not only disrupts semantic continuity but also restricts the optimization process, preventing embeddings from approaching the optimal solution.
Table 3: Image Similarity, Textual Alignment (BERTScore) and Prompt Interpretability comparison.
Dataset Method Image Similarity Textual Alignment Prompt Interpretability
@@
MS COCO PEZ 0.745 0.477 0.791 0.853 0.819 8,837.797
PH2P 0.789 0.467 0.801 0.851 0.823 11,078.994
VGD 0.725 0.509 0.826 0.885 0.853 665.418
EDITOR (Ours) 0.796 0.414 0.900 0.920 0.908 80.659
LAION PEZ 0.756 0.455 0.783 0.812 0.797 8,040.401
PH2P 0.826 0.427 0.794 0.810 0.802 9,621.671
VGD 0.769 0.466 0.800 0.818 0.808 700.754
EDITOR (Ours) 0.826 0.401 0.841 0.824 0.832 187.299
Flickr PEZ 0.722 0.442 0.774 0.857 0.811 9,425.353
PH2P 0.755 0.439 0.798 0.855 0.824 11,804.100
VGD 0.715 0.479 0.826 0.885 0.852 536.712
EDITOR (Ours) 0.776 0.424 0.892 0.915 0.900 90.320